mistyrose:

Solve the system of equations using the linear combination method. {5m+3n=41 3m−6n=9 Enter your answers in the boxes. m = n =

Shadow:

Do you know how to do linear combination?

mistyrose:

I have been stuck on this equation and im not really good at math

Shadow:

Linear combination is basically where you have two or more equations in which you need to add them together in order to solve for a variable.

Shadow:

The first step is to setup the equations so that when you add them together, there is only on variable remaining. This requires a simple look to see where and if you need to adjust the equations. Since no variables cancel out on their own through just adding them, I need to multiply one of the equations by some factor. \[5m+3n=41\] \[3m−6n=9\] It would seem the simplest way is to have 6n - 6n in order to get just have the variable m at the end. So we do: \[2(5m + 3n = 41) \] \[10m + 6n = 82\] Now we add \[10m + 6n = 82\] With \[3m - 6n = 9\] We get: \[13m = 91\] Divide by 13 \[m = 7\]

Shadow:

Does that make sense to you?

mistyrose:

Yes it does thank you

Shadow:

What do you think we do next?

mistyrose:

We have to figure out what N is?

Shadow:

Correct. And how do we go about doing that?

mistyrose:

How did you get 13 and 91?

Shadow:

10m + 6n = 82 3m - 6n = 9 Add these two equations together Like terms: 10m + 3m = 13m 6n - 6n = 0 82 + 9 = 91

mistyrose:

Oh I get it now

Shadow:

So do you know what we need to do next?

mistyrose:

Not gonna lie im not sure

Shadow:

It is a process known as substitution. Essentially you want to substitute the value of 7 for m into one of the variables. Then there is only one unknown, n, thus you can solve for it.

mistyrose:

okay and as we are talking and solving the equation im writing it down in my notes

Shadow:

What equation do you want to use?

Shadow:

@mistyrose

mistyrose:

sorry i was doing something for mom

Shadow:

</3

mistyrose:

I have never been smart when it came to any kind math

Shadow:

Review what we went over, them come back here.

Shadow:

Substitution: \[3m−6n=9\] \[3(7) - 6n = 9\] \[21 - 6n = 9\] Subtract 21 from both sides \[-6n = -12\] Divide both sides by -6 \[n = 2\]

mistyrose:

okay I will and thanks for the help

Shadow:

No problem